OpenStudy (aye_my_peeps):

cara walked 6 1/2 miles in 2 hrs. how long will it take him to walk 16 1/4 at the same speed? __hrs

OpenStudy (aye_my_peeps):

@theghostlycreeper

OpenStudy (anonymous):

1. Find his rate in miles/hour. 2. Rate * Time = distance so Time = Distance / Rate

OpenStudy (theghostlycreeper):

well my guss is just divide it by 4

OpenStudy (aye_my_peeps):

i still dont get it

OpenStudy (aye_my_peeps):

:(

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

\(\begin{array}{ccllll} miles&hrs \\\hline\\ 6\frac{1}{2}&2\\ 16\frac{1}{4}&x \end{array}\implies \cfrac{6\frac{1}{2}}{16\frac{1}{4}}=\frac{2}{x}\impliedby \textit{solve for "x"}\)